OpenStudy (anonymous):

Lesson 2: Multiplying and Dividing Radical Expressions ALG 2B Unit 1: Radical Functions and Rational Exponents (CONNECTIONS ACADEMY)

OpenStudy (preetha):

@SK_CA

OpenStudy (anonymous):

\[\sqrt[3]{4}\times \sqrt[3]{2000}\]

OpenStudy (anonymous):

would anyone like to explain to me how to do this

OpenStudy (displayerror):

Rather than radicals, make them fractions (ie. cube root = 1/3) \[4=2^2\] and \[2000=2*1000\] \[(2^2)^{\frac{1}{3}} * (2*1000)^{\frac{1}{3}}\] Recall that when you have something like \[(a^b)^c\] you multiply the exponents, so our equation becomes \[2^{\frac{2}{3}} * 2^{\frac{1}{3}} * 1000^{\frac{1}{3}}\] \[2^{\frac{2}{3} + \frac{1}{3}} * 10\] \[2*10=20\] So 20 is your answer.